The diameter of Earth is about twice that of Mars. The diameter of Mars is about three times that of Pluto. Together, the sum of the diameters of Earth, Mars, and Pluto is about 30,000 kilometers less than the diameter of Uranus. If the diameter of Uranus is about 52,000 kilometers, what equation would be used to find the approximate diameter of Pluto?
Answer by ankor@dixie-net.com(12692)

:Write an expression for each statement:
:
The diameter of Earth is about twice that of Mars.
E = 2M
:
The diameter of Mars is about three times that of Pluto.
M = 3P
or
P = (1/3)M or .33M
:
Together, the sum of the diameters of Earth, Mars, and Pluto is about 30,000 kilometers less than the diameter of Uranus.
E + M + P = U - 30000
:
Substitute for E and P:
2M + M + .33M = U - 30000
3.33M = U - 30000
:
If the diameter of Uranus is about 52,000 kilometers,
3.33M = 52000 - 30000
3.33M = 22000
M = 22000/3.33
M = 6600 km rounded off
what equation would be used to find the approximate diameter of Pluto?
:
P = (1/3)M
P = (1/3)6600
P = 2200 km
:
That's pretty close to it's diameter which is generally given as 2274 km
0 solutions